# Improving Plasticity in Online Continual Learning via Collaborative Learning

### v1

@@ Abstract
Online Continual Learning (CL) solves the problem of learning the ever-emerging new classification tasks from a continuous data stream. Unlike its offline counterpart, in online CL, the training data can only be seen once. Most existing online CL research regards catastrophic forgetting (i.e., model stability) as almost the only challenge. In this paper, we argue that the model’s capability to acquire new knowledge (i.e., model plasticity) is another challenge in online CL. While replay-based strategies have been shown to be effective in alleviating catastrophic forgetting, there is a notable gap in research attention toward improving model plasticity. To this end, we propose Collaborative Continual Learning (CCL), a collaborative learning based strategy to improve the model’s capability in acquiring new concepts. Additionally, we introduce Distillation Chain (DC), a novel collaborative learning scheme to boost the training of the models. We adapted CCL-DC to existing representative online CL works. Extensive experiments demonstrate that even if the learners are well-trained with state-of-the-art online CL methods, our strategy can still improve model plasticity dramatically, and thereby improve the overall performance by a large margin.
@@ 1 Introduction
Continual Learning (CL) [11, 35, 47, 14] aims to learn a sequence of tasks incrementally and encourage the neural network to gain more performance on the tasks at hand, without forgetting heretofore learned knowledge. CL can be done in two different manners [47, 4]: offline continual learning and online continual learning. In offline CL, the learner can have infinite access to all the training data of the current task it trains on and may go through the data for any epoch. Contrary to offline CL, in online CL, the training data for each task also comes continually in a data stream, and the learner can only see the training data once. Apart from the learning manner, there are also three different CL scenarios [25, 45, 32]: Task-incremental Learning (TIL), Domain-incremental Learning (DIL), and Class-incremental learning (CIL). In this paper, we focus on the CIL setting in online CL.
@@
Various online CL methods [38, 7, 6, 21, 22, 48] have been proposed to help the models learn continually on one-epoch data stream, with alleviated forgetting. Among them, replay-based methods have shown remarkable success, and current state-of-the-art methods rely heavily on memory replay to mitigate catastrophic forgetting [33, 19]. However, while most existing online CL research almost only focuses on improving model stability (i.e., alleviating catastrophic forgetting) in pursuit of better overall accuracy, the importance of model plasticity (i.e., the capability to acquire new knowledge) is greatly overlooked. Contrary to offline CL, where it is possible to gain high plasticity by iterating several epochs on the current task before proceeding to the subsequent task, the plasticity in online CL is more important because the training data can only be seen once. As shown in Fig. 1, compared to learning without memory replay, the replay-based methods implicitly alleviate the low plasticity issue to some extent. Also, it is possible to improve the plasticity with multiple updates trick on incoming samples [3]. However, the combination of memory replay and multiple updates does not bridge the plasticity gap, and multiple updates trick will also lead to higher catastrophic forgetting. Overall, the plasticity gap hinders the performance of online CL methods.
@@
In this paper, we claim that besides stability, the model’s ability to acquire new knowledge (i.e., model plasticity) is also vital in order to have a good overall accuracy. To shed light on how model plasticity and stability will impact the overall performance, we propose a quantitative link between plasticity, stability, and final accuracy, showing that both plasticity and stability play crucial roles in the overall performance.
@@
Guided by the quantitative relationship, we focus ourselves on the former-overlooked plasticity perspective. Inspired by the ability of collaborative learning to accelerate the convergence in non-continual scenarios [5], we incorporated collaborative learning in online CL and observed a similar phenomenon. To this end, we propose Collaborative Continual Learning with Distillation Chain (CCL-DC), a collaborative learning scheme that can be adapted to existing online CL methods. CCL-DC comprises two key components: Collaborative Continual Learning (CCL) and Distillation Chain (DC).
@@
CCL involves two peer continual learners to learn from the data stream simultaneously in a peer teaching manner, and it enables us to have more parallelism in optimization and provides more maneuverability to the continual learners. To the best of our knowledge, CCL is the first to involve collaborative learning techniques in online CL research. Moreover, to fully exploit the potential of collaborative learning in online CL scenarios, we proposed DC, an entropy regularization based optimization strategy explicitly designed for online CL.
@@
    We identify two important challenges in training online continual learners: plasticity and stability. Moreover, we propose a quantitative link between plasticity, stability, and final performance. Based on this, we find that plasticity is an important obstacle in online CL, which was greatly overlooked in the previous research;
@@
    To overcome the plasticity issue, we propose CCL-DC, a collaborative learning based strategy that can be seamlessly integrated into the existing methods and improve their performance by enhancing plasticity;
@@
    Extensive experiments show that CCL-DC can enhance the performance of existing methods by a large margin.
Figure 1: The plasticity (learning accuracy) and stability (relative forgetting, our metric proposed in Sec. 3) comparison of ER under different settings on CIFAR-100. For experiments with memory replay, the size of the memory buffer is set to 2,000. We can witness a plasticity gap between offline CL and online CL, even with memory replay and multiple update trick (memory iteration $>$ 1).
Figure 2: Overview of the proposed CCL-DC framework applied to a baseline online CL method. The proposed CCL-DC framework has two main components. The first one is CCL, which involves two peer continual learners that simultaneously learn from the data stream in a peer teaching manner. The second component is DC, which generates a chain of samples with varying levels of difficulty and feeds them to models to obtain a chain of logit distribution of different confidence levels. Then, in a collaborative learning approach, DC conducts distillation from less confident predictions to more confident predictions, to serve as a learned entropy regularization.
@@ 2 Related Work
Continual Learning methods can be classified into three different categories: regularization-based methods, parameter-isolation-based methods, and replay-based methods. Regularization-based methods [9, 29, 2, 26, 50] add extra regularization terms to balance the old and new tasks. Parameter-isolation-based methods [17, 40, 41, 39, 1] solve the problem explicitly by dynamically allocating task-specific parameters. Replay-based methods [38, 7, 6, 21, 22, 48, 34, 10, 15] maintain a small memory buffer that stores a few old training samples.
@@
These replay-based methods propose different strategies for alleviating catastrophic forgetting and improving the model stability. However, the importance of the model plasticity is greatly neglected in their research, despite their success in terms of final performance. In our work, these methods serve as the baselines and we adapted our strategy to these baselines to show the efficiency of our proposed approach.
@@
Collaborative learning [51, 20, 5, 52, 42] orients from online knowledge distillation (KD). Different from the conventional KD methods, online KD trains a cohort of deep networks from scratch in a peer-teaching manner. During the training process, the model imitates their peers and guides the training of other models simultaneously. DML [51] suggests peer student models learn from each other through the logit distillation between the probability distributions. Codistillation [5] is similar to DML and suggests the ensemble of peer networks can further improve the performance. More importantly, Codistillation shows that online KD can help the model converge faster on non-continual scenarios.
Despite the success of collaborative learning in non-continual scenarios, due to the lack of focus on plasticity, the research on collaborative learning in CL is still limited. To the best of our knowledge, there is no existing research using the collaborative learning technique to boost the training of online CL. Moreover, in our work, we propose DC, an entropy regularization based optimization strategy, which is designed to exploit the full potential of collaborative learning in online CL scenarios.
@@ 3 Plasticity and Stability in online CL
In this section, we revise the metric for model plasticity and propose a novel metric for model stability. In addition, we quantitatively derived the impact of model plasticity and stability on the final performance.
@@ 3.1 Model Plasticity
The model plasticity measures the learner’s capability to learn new knowledge when a new task arrives. Several different metrics have been proposed to measure the model plasticity [47, 9, 30, 37]. In our work, we evaluate the model plasticity with Learning Accuracy (LA) [37]. Formally, the Learning Accuracy for the $j$-th task is defined as:
@@
where $a_{j}^{i}$ is the accuracy evaluated on the test set of task $j$ after training the network from task 1 to task $i$. For an overall metric normalized against all tasks, the averaged Learning Accuracy is written as $LA=\frac{1}{T}\sum_{j=1}^{T}l_{j}$, and $T$ is the number of tasks in total.
@@ 3.2 Model Stability
The stability measures how much the model forgets given its current state. The most commonly used metric in previous CL research is the Forgetting Measure (FM) [9]. Intuitively, FM for the $j$-th task $fm_{j}^{k}$ reveals how much performance the model loses on a given task $j$, after training on task $k$, compared with its maximum performance obtained in the past:
$fm_{j}^{k}=\max_{i\in\{1,...,k-1\}}(a_{j}^{i}-a_{j}^{k}),\forall j<k.$ (2)
@@
In our work, instead of using FM as the stability metric, we propose forgetting measure on a relative basis, which we call Relative Forgetting (RF). There are two reasons for shifting from absolute forgetting to relative forgetting:
  1. 1.
    RF is more fair for methods with higher plasticity. This is because methods with poor plasticity will never have a large FM, and FM is capped by the maximum performance obtained by the learner in the past. Moreover, even if two learners lose the same absolute performance, the more plastic learner can still be regarded as forgetting less, because it has a higher peak performance and loses less proportion of its performance;
  2. 2.
    RF helps quantitatively derive the relationship between the model stability and final performance.
Intuitively, RF measures how much proportion of performance the model forgets. And RF for the $j$-th task after training on task $k$, can be defined as:
$f_{j}^{k}=\max_{i\in\{1,...,k\}}\left(1-\frac{a_{j}^{k}}{a_{j}^{i}}\right),\forall j\leq k.$ (4)
@@ 3.3 Impact on the Overall Performance
The model’s final average accuracy can be calculated by:
@@
With the definition of our plasticity metric (LA) and stability metric (RF), we can easily find the relationship between learning accuracy, relative forgetting, and accuracy:
@@
where we take the equal sign when $a_{j}^{j}=\max_{i\in\{1,...,i\}}a_{j}^{i}$. When generalizing the class-wise final accuracy $a_{j}^{T}$ to the average accuracy (AA), we need to take the dot product of the LA vector $[l_{1},...,l_{T}]$ with RF vector $[f_{1}^{T},...,f_{T}^{T}]$ which is trivial. More intuitively, in practice, we can make the approximation with:
@@ 4 Proposed Method
In this section, we first justify our motivation with the findings in Sec. 3. Then, we introduce our proposed strategy: Collaborative Continual Learning and Distillation Chain. Finally, we show how to adapt our proposed strategy to the existing online CL methods and boost their plasticity.
@@ 4.1 Motivation Justification
Online continual learners aim to continuously adapt to non-stationary data streams, efficiently acquiring new knowledge while retaining previously learned information. In current online CL research, almost all of the efforts focus on alleviating catastrophic forgetting, and the importance of “learning capability” on new knowledge is greatly overlooked. However, our finding in Sec. 3 shows both plasticity and stability play important roles in achieving decent final performance.
While replay-based methods were originally designed to tackle forgetting issues, Fig. 1 demonstrates that they can implicitly mitigate the plasticity gap. Nonetheless, as shown in Fig. 1, the limited plasticity is still a significant barrier to the performance, even with replay. To this end, we explicitly focus on the plasticity perspective.
The potential of collaborative learning to improve convergence in non-continual scenarios [5] positions it as a promising candidate for enhancing plasticity. With the apparent lack of focus on plasticity, collaborative learning has yet to be leveraged to boost convergence of online continual learners. In our research, we propose to exploit collaborative learning convergence properties for improving plasticity. We find that similar to non-continual scenarios, collaborative learning strategy can boost convergence by allowing more parallelism in the training and more maneuverability of the continual learners. Moreover, to fully take advantage of collaborative learning, we also propose Distillation Chain (DC), an entropy regularization based optimization strategy in collaborative learning specifically designed for online CL.

### v2

# breaklines

@@ Abstract
Online Continual Learning (CL) solves the problem of learning the ever-emerging new classification tasks from a continuous data stream. Unlike its offline counterpart, in online CL, the training data can only be seen once. Most existing online CL research regards catastrophic forgetting (i.e., model stability) as almost the only challenge. In this paper, we argue that the model’s capability to acquire new knowledge (i.e., model plasticity) is another challenge in online CL. While replay-based strategies have been shown to be effective in alleviating catastrophic forgetting, there is a notable gap in research attention toward improving model plasticity. To this end, we propose Collaborative Continual Learning (CCL), a collaborative learning based strategy to improve the model’s capability in acquiring new concepts. Additionally, we introduce Distillation Chain (DC), a collaborative learning scheme to boost the training of the models. We adapt CCL-DC to existing representative online CL works. Extensive experiments demonstrate that even if the learners are well-trained with state-of-the-art online CL methods, our strategy can still improve model plasticity dramatically, and thereby improve the overall performance by a large margin. The source code of our work is available at https://github.com/maorong-wang/CCL-DC.
@@ 1 Introduction
Continual Learning (CL) [10, 32, 43, 13] aims to incrementally learn a sequence of tasks and enhance the neural network’s performance on the current tasks, without forgetting previously learned knowledge. CL can be done in two different manners [43, 4]: offline and online. In offline CL, the learner can have infinite access to all the training data of the current task that it trains on and may go through the data for any epoch. Contrary to offline CL, in online CL, the training data for each task also comes continually in a data stream, and the learner can only see the training data once. Apart from the learning approach, there are also three different CL scenarios [23, 42, 29]: Task-Incremental Learning (TIL), Domain-Incremental Learning (DIL), and Class-Incremental learning (CIL). In this paper, we focus on the CIL setting in online CL.
@@
Various online CL methods [35, 7, 6, 19, 20, 44, 31] have been proposed to help the models learn continually on one-epoch data stream, with alleviated forgetting. Among them, replay-based methods have shown remarkable success, and current state-of-the-art methods rely heavily on memory replay to mitigate catastrophic forgetting [30, 17]. However, while most existing online CL research primarily only focuses on improving model stability (i.e., alleviating catastrophic forgetting) to achieve better overall accuracy, the importance of model plasticity (i.e., the capability to acquire new knowledge) is often overlooked. While in offline CL, it is possible to gain high plasticity by iterating several epochs on the current task before proceeding to the subsequent task, acquiring plasticity in online CL tends to be more challenging because the training data is only available once. As shown in Fig. 1, compared to learning without memory replay, the replay-based methods implicitly alleviate the low plasticity issue to some extent. Also, plasticity can be improved by employing a technique involving multiple updates on incoming samples [3]. However, the combination of memory replay and multiple updates fails to bridge the existing plasticity gap between online and offline CL, and the use of multiple updates can lead to increased catastrophic forgetting. Overall, the plasticity gap adversely affects the performance of online CL methods.
@@
In this paper, we argue that plasticity is particularly crucial and challenging to acquire in online CL, and even more so than in offline CL. Thus, we shed light on how model plasticity and stability will impact the overall performance. Furthermore, we propose a quantitative link between plasticity, stability, and final accuracy, showing that the plasticity gap between offline and online must be reduced to improve overall performance.
@@
Guided by the quantitative relationship, we focus ourselves on the previously overlooked plasticity perspective. Inspired by the ability of collaborative learning to accelerate the convergence in non-continual scenarios [5], we have incorporated collaborative learning in online CL and observed a similar phenomenon. Consequently, we propose the Collaborative Continual Learning with Distillation Chain (CCL-DC), a collaborative learning scheme that can be adapted to existing online CL methods. CCL-DC consists of two key components: Collaborative Continual Learning (CCL) and Distillation Chain (DC).
@@
CCL involves two peer continual learners to learn from the data stream simultaneously in a peer teaching manner, enhancing optimization parallelism and offering greater flexibility to the continual learners. To the best of our knowledge, CCL is the first method to incorporate collaborative learning techniques in online CL research. Moreover, to fully exploit the potential of collaborative learning in online CL scenarios, we propose DC, an entropy regularization based optimization strategy, specifically designed for online CL.
@@
    We establish a quantitative link between plasticity, stability, and final performance. Based on this, we demonstrate that plasticity is an important obstacle in online CL, which was greatly overlooked in the previous research;
@@
    To overcome the plasticity issue, we introduce CCL-DC, a collaborative learning based strategy that can be seamlessly integrated into the existing methods, significantly improving their performance by boosting plasticity;
@@
    Comprehensive experiments show that CCL-DC enhances the performance of existing methods by a large margin.
Figure 1: The comparison of plasticity (learning accuracy) and stability (relative forgetting, our metric proposed in Sec. 3) of Experience Replay (ER) [35] under varisous settings on CIFAR-100. For experiments with memory replay, the size of the memory buffer is set to 2,000. We can witness a plasticity gap between offline CL and online CL, even with memory replay and multiple update trick (memory iteration $>$ 1).
Figure 2: Overview of the proposed CCL-DC framework applied to a baseline online CL method. The proposed CCL-DC framework has two main components. The first one is CCL, which involves two peer continual learners that simultaneously learn from the data stream in a peer teaching manner. The second component, DC, generates a chain of samples with varying levels of difficulty and feeds them into models to produce a chain of logit distribution of different confidence levels. Then, in a collaborative learning approach, DC conducts distillation from less confident predictions to more confident predictions, to serve as a form of learned entropy regularization.
@@ 2 Related Work
Continual learning can be classified into three different categories: regularization-based, parameter-isolation-based, and replay-based. Regularization-based methods [8, 27, 2, 24, 46] add extra regularization terms to balance the old and new tasks. Parameter-isolation-based methods [16, 37, 38, 36, 1] solve the problem explicitly by dynamically allocating task-specific parameters. Replay-based methods [35, 7, 6, 19, 20, 44, 31, 9, 14] maintain a small memory buffer that stores a few old training samples.
@@
These replay-based methods propose different strategies for alleviating catastrophic forgetting and improving the model stability. However, the importance of the model plasticity is greatly neglected in their research, despite their success in terms of final performance. In our work, these methods serve as the baselines and we adapted our strategy to these baselines to show the effectiveness of our proposed approach.
@@
Collaborative learning [47, 18, 5, 48, 39] originates from online knowledge distillation (KD). Different from the conventional KD methods, online KD involves training a cohort of deep networks from scratch in a peer-teaching manner. During the training process, the models imitate their peers and guides the training of other models simultaneously. Deep Mutual Learning (DML) [47] suggests peer student models learn from each other through the logit distillation between the probability distributions. Codistillation [5] is similar to DML and suggests the ensemble of peer networks can further improve the performance. More importantly, Codistillation shows that online KD can help the model converge faster on non-continual scenarios.
Despite the success of collaborative learning in non-continual scenarios, due to the lack of focus on plasticity, the research on collaborative learning in CL is still limited. To the best of our knowledge, there is no existing research using the collaborative learning technique to boost the training of online CL. Moreover, in our work, we introduce the DC, an entropy regularization based optimization strategy, specifically tailored to fully leverage the potential of collaborative learning in online CL scenarios.
@@ 3 Plasticity and Stability in online CL
In this section, we revise the metric for model plasticity and propose a novel metric for model stability. In addition, we quantitatively analyze the impact of model plasticity and stability on the final performance.
@@ 3.1 Model Plasticity
The model plasticity quantifies the learner’s ability to learn new knowledge when a new task arrives. Several different metrics have been proposed to measure the model plasticity [43, 8, 28, 34]. In our work, we evaluate the model plasticity with Learning Accuracy (LA) [34]. Formally, the LA for the $j$-th task is defined as:
@@
where $a_{j}^{i}$ is the accuracy evaluated on the test set of task $j$ after training the network from task 1 to task $i$. For an overall metric normalized against all tasks, the averaged LA is written as $LA=\frac{1}{T}\sum_{j=1}^{T}l_{j}$, and $T$ is the number of tasks in total.
@@ 3.2 Model Stability
The stability assesses the extent of knowledge retention or loss within a model in its present state. The most commonly used metric in previous CL research is the Forgetting Measure (FM) [8]. Intuitively, FM for the $j$-th task, denoted as $fm_{j}^{k}$, quantifies the decline in performance on a given task $j$, after the model is trained on task $k$, relative to its highest past performance on task $j$:
$fm_{j}^{k}=\max_{i\in\{1,...,k-1\}}(a_{j}^{i}-a_{j}^{k}),~{}\forall j<k.$ (2)
@@
In our work, we propose a novel stability metric, termed Relative Forgetting (RF), as an alternative to the traditional Forgetting Measure (FM). RF measures how much proportion of performance the model forgets. Specifically, RF for the $j$-th task after training on task $k$ is defined as:
$f_{j}^{k}=\max_{i\in\{1,...,k\}}\left(1-\frac{a_{j}^{k}}{a_{j}^{i}}\right),~{}\forall j\leq k.$ (4)
@@
$a_{j}^{i}$ $\mathcal{T}_{1}$ $\mathcal{T}_{2}$ $\mathcal{T}_{3}$ $\mathcal{T}_{4}$ $\mathcal{T}_{5}$ $FM_{k}$ $RF_{k}$
$\mathcal{T}_{1}$ ${\color[rgb]{0,.5,.5}30}/{\color[rgb]{.75,.5,.25}15}$
$\mathcal{T}_{2}$ ${\color[rgb]{0,.5,.5}25}/{\color[rgb]{.75,.5,.25}12.5}$ ${\color[rgb]{0,.5,.5}25}/{\color[rgb]{.75,.5,.25}12.5}$ ${\color[rgb]{0,.5,.5}5}/{\color[rgb]{.75,.5,.25}2.5}$ $/{\color[rgb]{.75,.5,.25}8.33}$ 8.33
$\mathcal{T}_{3}$ ${\color[rgb]{0,.5,.5}20}/{\color[rgb]{.75,.5,.25}10}$ ${\color[rgb]{0,.5,.5}20}/{\color[rgb]{.75,.5,.25}10}$ ${\color[rgb]{0,.5,.5}20}/{\color[rgb]{.75,.5,.25}10}$ ${\color[rgb]{0,.5,.5}7.5}/{\color[rgb]{.75,.5,.25}3.75}$ $/{\color[rgb]{.75,.5,.25}17.78}$ 17.78
$\mathcal{T}_{4}$ ${\color[rgb]{0,.5,.5}15}/{\color[rgb]{.75,.5,.25}7.5}$ ${\color[rgb]{0,.5,.5}15}/{\color[rgb]{.75,.5,.25}7.5}$ ${\color[rgb]{0,.5,.5}15}/{\color[rgb]{.75,.5,.25}7.5}$ ${\color[rgb]{0,.5,.5}15}/{\color[rgb]{.75,.5,.25}7.5}$ ${\color[rgb]{0,.5,.5}10}/{\color[rgb]{.75,.5,.25}5}$ $/{\color[rgb]{.75,.5,.25}28.75}$ 28.75
$\mathcal{T}_{5}$ ${\color[rgb]{0,.5,.5}10}/{\color[rgb]{.75,.5,.25}5}$ ${\color[rgb]{0,.5,.5}10}/{\color[rgb]{.75,.5,.25}5}$ ${\color[rgb]{0,.5,.5}10}/{\color[rgb]{.75,.5,.25}5}$ ${\color[rgb]{0,.5,.5}10}/{\color[rgb]{.75,.5,.25}5}$ ${\color[rgb]{0,.5,.5}10}/{\color[rgb]{.75,.5,.25}5}$ ${\color[rgb]{0,.5,.5}12.5}/{\color[rgb]{.75,.5,.25}6.25}$ $/{\color[rgb]{.75,.5,.25}42}$ 42
Table 1: Forgetting Measure (%) and Relative Forgetting (%) of two example learners (Learner 1 in teal and Learner 2 in brown) in a continual setting of five tasks with two classes per task. Each element $a_{j}^{i}$ at row $j$ and column $i$ is the per task accuracy evaluated on the test set of task $j$ after training the network from task 1 to task $i$.
There are two advantages for shifting from absolute forgetting to relative forgetting:
  1. 1.
    RF is fairer for methods with higher plasticity. As shown in Table 1, when comparing two continual learners, one with high plasticity and the other with low, using both forgetting metrics, it is observed that despite having identical RF values, the model with lower plasticity shows a lower FM. This indicates that the FM metric might unfairly favor models with poorer initial performance, even when the relative decline in performance is equivalent across models;
  2. 2.
    RF facilitates a quantitative analysis of the relationship between model stability and final performance. This enables a more nuanced understanding of how stability influences overall model efficacy, thereby guiding improvements in continual learning strategies.
@@ 3.3 Impact on the Overall Performance
The model’s final average accuracy can be calculated as:
@@
With the establishment of our plasticity metric (LA) and stability metric (RF), we can easily find the relationship between learning accuracy, relative forgetting, and accuracy:
@@
assuming equality when $a_{j}^{j}=\max_{i\in\{1,...,j\}}a_{j}^{i}$. When extending the class-specific final accuracy $a_{j}^{T}$ to the final AA, we consider the dot product of the LA vector $[l_{1},...,l_{T}]$ with the RF vector $[f_{1}^{T},...,f_{T}^{T}]$, which simplifies the calculation. More practically, this relationship can be approximated with:
@@ 4 Proposed Method
In this section, we first justify our motivations. Then, we introduce our proposed strategy: Collaborative Continual Learning and Distillation Chain. Finally, we show how to adapt our proposed strategy to the existing online CL methods and boost their plasticity.
@@ 4.1 Motivation Justification
Figure 3: Conceptual diagram of the training framework, when distilling from an untrained network $\theta^{2}$ to suppress the confidence of network $\theta^{1}$.
Dataset CIFAR-100
Memory Size $M$ M=1000 M=2000 M=5000
ER 24.47±0.72 31.89±1.45 39.41±1.81
ER + Untrained Distillation 27.07±1.20 34.84±0.64 41.15±1.16
Table 2: The performance of network $\theta^{1}$ when distilling from untrained network $\theta^{2}$ on CIFAR-100. All numbers are average over 10 runs.
Plasticity matters in Online CL.
Online continual learners are designed to continuously adapt to non-stationary data streams, efficiently acquiring new knowledge while retaining previously learned information. In current online CL research, the emphasis has predominantly been on mitigating catastrophic forgetting. Even among state-of-the-art methods such as [19, 44], plasticity is often compromised in the quest for improved stability. However, our finding in Sec. 3 shows that both plasticity and stability are crucial for achieving satisfactory final performance, with plasticity proving particularly difficult to achieve in online CL settings. To this end, we deliberately prioritize the plasticity aspect.
The potential of collaborative learning to improve convergence in non-continual scenarios [5] positions it as a promising candidate for enhancing plasticity. With the apparent lack of focus on plasticity, collaborative learning has yet to be leveraged to boost convergence of online continual learners. In our research, we propose to exploit collaborative learning convergence properties for improving plasticity. We find that similar to non-continual scenarios, collaborative learning strategy can boost convergence by allowing more parallelism in the training and more maneuverability of the continual learners.
Overconfidence hurts Online CL.
In conventional supervised learning, it is well established that excessive confidence can harm the generalization ability. To tackle this issue, many methods such as label smoothing [41], knowledge distillation [45], and confidence penalty [33] are proposed.
In online CL, we find a counter-intuitive, yet similar phenomenon. As shown in Fig. 3, while we train network $\theta^{1}$ with the classification loss, we initiate another network $\theta^{2}$ at the beginning of the training. During the continual training, we stagnate the network $\theta^{2}$ (i.e., Kaiming initialize and stay untrained) and train the network $\theta^{1}$ with both classification loss and the distillation loss (Kullback-Leibler divergence loss) from untrained network $\theta^{2}$. The experimental result in Table 2 shows a decent performance gain compared with independent training. The distillation from the prediction of untrained network $\theta^{2}$ to network $\theta^{1}$ serves as a regularization to suppress the overall confidence, and the performance gain under this condition makes it evident that, in continual scenarios, overconfidence will also harm the performance.
To tackle the overconfidence problem, we propose Distillation Chain (DC), an entropy regularizer to suppress the overall confidence level.